OpenStudy (anonymous):

Solve. Assume the exercise describes a linear relationship. When making a telephone call using a calling card, a call lasting 3 minutes costs $1.05. A call lasting 12 minutes costs $2.85. Let y be the cost of making a call lasting x minutes using a calling card. Write a linear equation that models the cost of making a call lasting x minutes. Choose one answer. a. y = -0.2x + 1.65 b. y = 0.2x - 9.15 c. y = 0.2x + 0.45 d. y = 5x - 279/20

OpenStudy (dumbcow):

given 2 points: (3,1.05) and (12,2.85) the slope is: m = (2.85-1.05)/(12-3) = 1.8/9 = .2 y-1.05 = .2(x-3) y = .2x +.45